### v2

# CEHR-XGPT: A Scalable Multi-Task Foundation Model for Electronic Health Records

Abstract
Electronic Health Records (EHRs) provide a rich, longitudinal view of patient health and hold significant potential for advancing clinical decision support, risk prediction, and data-driven healthcare research. However, most artificial intelligence (AI) models for EHRs are designed for narrow, single-purpose tasks, limiting their generalizability and utility in real-world settings. Here, we present CEHR-XGPT, a general-purpose foundation model for EHR data that unifies three essential capabilities - feature representation, zero-shot prediction, and synthetic data generation - within a single architecture. To support temporal reasoning over clinical sequences, CEHR-XGPT incorporates a novel time-token-based learning framework that explicitly encodes patients’ dynamic timelines into the model structure. CEHR-XGPT demonstrates strong performance across all three tasks and generalizes effectively to external datasets through vocabulary expansion and fine-tuning. Its versatility enables rapid model development, cohort discovery, and patient outcome forecasting without the need for task-specific retraining.
Keywords
EHR Foundation Model, Synthetic Electronic Health Records, Patient Representation, Observational Medical Outcomes Partnership - Common Data Model, Observational Health Data Sciences and Informatics
1 Introduction
Electronic health records (EHRs) capture rich, longitudinal information about patients across diverse clinical settings. This wealth of data presents significant opportunities to build generalizable and clinically useful AI models. However, it is highly challenging to model such data reliably. EHRs differ from typical sequential data in several important ways: they are irregularly sampled, temporally sparse, and often contain heterogeneous modalities such as diagnoses, procedures, lab results, and medications. Accurately modeling the temporality and structure of EHR data is essential to capture the underlying progression of patient health and produce reliable, context-aware predictions [1, 2, 3].
With the advent of general-purpose foundation models, such as ChatGPT—trained on large, diverse datasets using self-supervised learning rather than single-task objectives—researchers have begun developing analogous foundation models for EHRs [4, 5]. These EHR foundation models aim to capture the full potential of longitudinal patient data by learning generalizable representations that can support a variety of downstream tasks with improved performance.
Recent EHR foundation models have been applied to specific applications such as disease prediction, patient outcome forecasting, synthetic data generation, extracting patient representations, and propensity score modeling for causal effect estimation [6, 7, 8, 9, 5, 10, 11, 12, 13, 14, 15, 16]. However, most existing models have been designed and evaluated with a narrow focus—typically excelling at one task type, such as representation learning or predictive modeling, but not both.
We present CEHR-XGPT (pronounced “seer-ex-gpt”), a unified EHR foundation model that supports three core capabilities within a single framework: feature representation, zero-shot prediction, and synthetic data generation. By integrating all three, our model demonstrates a broader and more flexible utility compared to prior approaches.
  • Feature Representation: The model generates patient embeddings from sequences of medical events, enabling a variety of downstream tasks such as disease prediction, patient clustering, and propensity score matching.
  • Zero-Shot Prediction: The model can predict future patient events directly from prompts, without requiring task-specific training or fine-tuning. This enables rapid evaluation of new prediction tasks in settings with limited labeled data.
  • Synthetic Data Generation: The model learns the sequence distribution of real patient data and generates synthetic timelines that preserve key statistical properties and inter-variable dependencies, supporting tasks like simulation, data sharing, and augmentation.
The versatility of CEHR-XGPT stems from a novel time-token–based learning framework that explicitly encodes patients’ dynamic timelines into the model architecture, preserving the full temporal structure and sequence dependencies inherent in longitudinal EHR data. To our knowledge, CEHR-XGPT is the first multi-task foundation model to simultaneously support all three core capabilities within a unified framework.
2 Related work
There has been a growing number of studies on EHR foundation models, each targeting slightly different capabilities, including feature representation, zero-shot prediction, and synthetic data generation.
Feature Representation: Recent works such as MOTOR, along with models like Llama and Mamba reported by Wornow et al., have focused on the feature representation capabilities of EHR foundation models using linear probing [7, 8, 5]. Linear probing is a model evaluation technique in which a lightweight classifier (typically logistic regression) is trained on fixed patient representations generated by a pretrained foundation model. This approach allows researchers to assess how well the model’s learned embeddings encode clinically relevant information, without further training for specific tasks. In the context of EHR applications, linear probing offers several advantages: it requires only a single forward pass through the model, is computationally efficient, and avoids the need for fine-tuning, making it attractive for rapid evaluation and deployment. In contrast, other models such as MED-BERT, BEHRT, TransformerEHR, and EHRMamba require additional fine-tuning for each downstream task after pretraining [11, 17, 10, 18, 13, 14, 19, 20, 21]. These models differ in two main aspects. First, they adopt various architectural choices, including encoder-only, decoder-only, and encoder–decoder designs. Second, they incorporate temporal information in different ways, such as through positional embeddings, age embeddings, time embeddings, or, more recently, discrete time tokens.
Zero-Shot Prediction: Several EHR Foundation Models, such as Foresight and ETHOS, built on causal language model architectures, focus on zero-shot prediction without any task-specific training [22, 15, 12]. In Foresight, the authors first used NLP tools to extract clinical concept codes from unstructured notes, which they combined with structured EHR data to construct full patient timelines. To encode demographic context, age tokens were inserted into the timeline, with a new token added whenever the patient’s age increased. Foresight achieved strong performance in forecasting future conditions, demonstrating high precision and recall across ranked predictions (e.g., top-1, top-5, top-10). A qualitative evaluation was also conducted, in which five clinicians assessed predictions generated from 34 synthetically constructed patient timelines. Foresight achieved a high relevance score, with 33 out of 34 predictions (95%) deemed clinically meaningful. Similarly, ETHOS, trained primarily on MIMIC-IV, demonstrated strong zero-shot capabilities, matching or surpassing prior SOTA models on several clinical prediction benchmarks. However, both models employed tokenization strategies that compromised temporal fidelity. In Foresight, age tokens served as coarse temporal markers, making it impossible to distinguish between events occurring 1 day apart versus 90 days apart if the patient’s age remained constant. In ETHOS, more granular time tokens (e.g., 6–12 hours, 3–6 months) were introduced to better model ICU timelines, but this approach still struggled to capture time gaps across discrete hospital admissions. While such models are well-suited for forecasting tasks that tolerate temporal compression, they are less appropriate for applications like synthetic data generation or fine-grained trajectory modeling, which require high temporal resolution.
Synthetic Data Generation: Structured EHRs are characterized as irregular time-series data with high-dimensional features at each time point. Most prior work has focused on deep learning-based methods without adequately preserving the temporal structure of clinical events [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 6, 19, 34]. For example, HALO, a state-of-the-art (SOTA) model for generating synthetic EHR data, incorporates inter-visit time intervals but fails to represent inpatient visit durations [6]. This omission makes it unsuitable for tasks that require accurate modeling of patient trajectories over time, such as identifying 30-day readmissions, thereby limiting its evaluation to simpler tasks like predicting the presence of medical codes rather than supporting more complex phenotype modeling. Moreover, HALO discretizes inter-visit intervals into coarse buckets and samples from a uniform distribution, which can distort realistic patient timelines. In real-world outpatient care, for instance, 7- and 14-day intervals are much more common. Finally, HALO was evaluated separately on inpatient and outpatient datasets rather than on a unified EHR dataset that spans both care settings, further limiting its generalizability.
These capabilities require varying levels of patient timeline preservation and generative capacity. For feature representation, full patient timelines are not strictly necessary—some foundation models are not generative and simply order medical events chronologically without modeling temporal dependencies [18, 10]. In zero-shot prediction, models are not required to reconstruct complete timelines; some degree of timeline shrinkage or abstraction can be tolerated without significantly impacting performance [22, 15]. In contrast, synthetic data generation requires high-fidelity modeling of entire patient trajectories. To generate realistic synthetic sequences, it is essential to preserve the full temporal structure and the sequence dependencies inherent in the original data [6, 35]. As a result, a strong synthetic data generator should, in theory, also perform well on zero-shot prediction and feature representation tasks; however, the reverse does not necessarily hold.
Figure 1: CEHR-XGPT is built on the GPT-2 architecture with a specialized patient representation designed to capture the full longitudinal trajectory of each patient. Two additional learning objectives were applied to artificial time token embeddings to improve the modeling of temporal information, namely, time decomposition and time-to-event prediction. CEHR-XGPT supports three key capabilities: patient-level feature representation, zero-shot prediction, and synthetic data generation.
3 Methods
3.1 Data Source and Preprocessing
The training patient sequences were derived from a subset of the Columbia University Irving Medical Center-New York Presbyterian Hospital OMOP database. This subset includes data on conditions, medications, and procedures from patient medical histories. Unknown concepts (i.e., $\texttt{concept\_id}=0$) were excluded from all domains except the visit type when constructing the patient sequences using our proposed representation. Additionally, we excluded patients with fewer than 20 tokens to ensure sufficient longitudinal information for model training. The final dataset included approximately 2.6 million patients for training and an additional 1 million patients reserved for evaluation. To enable comparison with other baselines such as MOTOR [36, 8], we also converted the CUIMC-NYP OMOP dataset into the Medical Event Data Standard (MEDS) format.
3.2 Representing Patient Trajectory
Each patient’s EHR record is modeled as a sequence of events, where the demographic information (comprising start_year, start_age, gender, and race) is positioned at the beginning. This is followed by the patient’s visits, arranged chronologically. Due to the irregular nature of patient trajectories, where time intervals between visits vary, we introduce artificial time tokens (ATT) to quantify these intervals in days. For intervals shorter than 1080 days, a corresponding ATT token, such as $D10$ for a 10-day gap, is used. Intervals exceeding 1080 days are denoted by a single [LT] token (long-term). Each visit is encapsulated between two special tokens, [VS] and [VE], to mark the start and end of a visit, respectively. Immediately following [VS], a visit-type token [VT] specifies the type of visit. For inpatient visits, a discharge-type token is inserted at the end of the visit right before [VE]. The structure of a patient sequence is thus represented as:
$P=\{$ $\texttt{[start\_year]},\texttt{[start\_age]},\texttt{[gender]},\texttt{[race]}$
$\texttt{[VS] [VT]},v_{1},\texttt{[VE]},ATT$
$\texttt{[VS] [VT]},v_{2},\texttt{[discharge\_type][VE]},ATT$
$\dots$
$\texttt{[VS] [VT]},v_{i},\texttt{[VE]}\}$
where $v_{i}$ represents an arbitrary visit comprising a chronologically ordered list of medical events (denoted as $c_{i}$), and $v_{i}=\{c_{i1},c_{i2},\dots,c_{ij}\}$.
3.3 Model Architecture
We developed our model architecture based on the GPT-2 framework [37], tailored to effectively capture the sequence distribution. To enhance the model’s efficiency, we excluded the positional embeddings from the model due to two primary considerations:
  1. 1.
    The temporal information is inherently captured through the use of ATT tokens, making separate positional embeddings unnecessary.
  2. 2.
    As demonstrated by Haviv et al [38], transformer-based language models such as GPT can learn and leverage positional information implicitly, even in the absence of explicit or structured positional encodings. The causal attention mechanism employed by GPT inherently enables the model to account for temporally ordered patient histories by conditioning on all prior tokens in a sequence, without requiring additional structured inputs such as positional embeddings.
Our preliminary experiments supported this design choice, showing that trainable positional embeddings across different positions tended to converge to similar vectors, suggesting their limited utility in our setup. Additionally, we incorporated two specialized learning objectives, Time Decomposition (TD) and Time to Event (TTE), designed to enhance the representation of time intervals within the model. The overall architecture of the CEHR-XGPT foundation model is illustrated in Figure 1.
Figure 2: The distribution of time tokens between visits is not uniform but instead biased toward specific intervals, such as 7 and 14 days.
The TD learning objective introduces a structural constraint on the ATTs to address the skewness in the distribution of time intervals, as depicted in Figure 2. While short-term intervals are more frequent, there is a noticeable long tail extending towards the long-term ATTs. The TD principle allows for the decomposition of any time interval into years, months, and days. For example, $D1$ would be decomposed into 0 year, 0 month, and $1$ day, whereas $D396$ would be translated to $1$ year, $1$ month, and $1$ day.
To implement this, we added a TD prediction head at the positions corresponding to ATT tokens within the output layer. Each ATT embedding is split into three sub-embeddings that predict the time components of year, month, and day. This multiclass classification scenario uses cross-entropy loss to compute the loss term, allowing the model to leverage shared TD predictions across different timescales to improve the accuracy of long-term ATT predictions. The TD loss term can be expressed as follows,
$[e^{year}_{ij}$ $,e^{month}_{ij},e^{day}_{ij}]=split(e^{att}_{ij})$
$\hat{y}^{year}_{ij}=W_{year}\times e^{year}_{ij}$
$\hat{y}^{month}_{ij}=W_{month}\times e^{month}_{ij}$
$\hat{y}^{day}_{ij}=W_{day}\times e^{day}_{ij}$
$TD_{ij}=$ $-\sum P(y^{year}_{ij})\log P(\hat{y}^{year}_{ij}\mid e_{ij}^{year})$
$-\sum P(y^{month}_{ij})\log P(\hat{y}^{month}_{ij}\mid e_{ij}^{month})$
$-\sum P(y^{day}_{ij})\log P(\hat{y}^{day}_{ij}\mid e_{ij}^{day})$
where $e^{att}_{ij}$ denotes the embedding representation of an ATT at the jthjth position of the ithith patient, $e^{year}_{ij}$, $e^{month}_{ij}$, $e^{day}_{ij}$ represent the sub-embeddings corresponding to year, month and day. $W_{year}$, $W_{month}$, $W_{day}$ represent the linear maps to transform the sub-embeddings to the corresponding time component predictions.
In addition to incorporating TD, we introduced a Time-To-Event (TTE) learning objective to enhance the semantic depth of ATTs. When ATT embeddings are trained solely through next-token prediction, driven primarily by co-occurrence statistics, the model is limited to learning only the relative positions of ATTs in the embedding space, reducing their semantic richness. To address this limiation, TTE uses a Gamma probability distribution, parameterized by a feed-forward layer, at positions corresponding to ATTs. The model is optimized by minimizing the negative log-likelihood of the actual time intervals under a Gamma distribution. The resulting TTE loss term is defined as:
$\alpha_{ij},\beta_{ij}=feed\_forward(e^{att}_{ij})$
$TTE_{ij}=-\log P(\Delta T_{ij};Gamma(\alpha_{ij},\beta_{ij}))$
where $e^{att}_{ij}$ denotes the embedding representation of an ATT at the jthjth position of the ithith patient, $\alpha_{ij}$ and $\beta_{ij}$ are the parameters for the Gamma distribution generated by a feed-forward network, $\Delta T_{ij}$ is the time difference in days represented by $e^{att}_{ij}$.
The overall loss function of CEHR-XGPT is defined below, which comprises the original next token prediction (NTP), denoted as $L^{ntp}_{ij}$, and the two additional learning objectives defined above,
$L=$ $\sum_{i}\sum_{j}\Big{(}L^{ntp}_{ij}+\mathbbm{1}[c_{ij}\in ATTs]\big{(}TTE_{ij}+TD_{ij}\big{)}\Big{)}$
No. of visits
No. of concepts
mean 16 117
std 19 278
min 2 8
25% 1 14
50% 3 32
75% 11 98
99% 216 1775
max 1560 9994
Table 1: Summary statistics of the CUIMC-NYP OMOP training data
3.4 Training Setup
The model uses a 4096-token context window, 16 transformer decoder blocks, 12 attention heads, and 768 dimensional embeddings and hidden units. A dropout rate of $0.1$ was applied to reduce overfitting. Input sequences exceeding 4096 tokens were truncated to fit the context window. Summary statistics of the training data are provided in Table 1, and a full model configuration is listed in Supplementary Table 7. The model was trained for 10 epochs on a single NVIDIA A6000 GPU using sample packing, with a batch size of up to 16,384 tokens and an initial learning rate of 0.001. To enable recovery and intermediate evaluation without retraining from scratch, checkpoints were saved every 20,000 training steps (approximately per epoch).
3.5 Time tokens V.S. Time embeddings
In this section, we provide a theoretical justification for inserting time tokens into patient sequences, which may yield better results than the traditional summation strategy of adding time embeddings to concept embeddings - a common approach in prior work [18, 17, 14].
To illustrate the limitations of the summation approach, we consider a simple logical function where the output depends on the time interval between two binary events $x_{1},x_{2}\in\{0,1\}$, with timestamps $t_{1},t_{2}$, where $t_{1}\leq t_{2}$:
  1. 1.
    If $t_{2}-t_{1}\leq 7$, then $y=\text{XOR}(x_{1},x_{2})$.
  2. 2.
    If $t_{2}-t_{1}>7$, then $y=\text{AND}(x_{1},x_{2})$.
When using a token-based input like $\vec{x}_{\text{token}}^{T}=[x_{1},\Delta t,x_{2}]$, a simple feedforward neural network can be constructed to correctly route the computation through the appropriate logic gate based on the time interval $\Delta t$. We demonstrate this using an example input of $\vec{x}_{\text{token}}^{T}=[0,6,1]$, along with hand-designed weights and thresholds, showing that the model can successfully activate XOR or AND logic depending on the interval.
We define matrices, $W$, and bias vectors, $\beta^{T}$:
$W_{1}=\begin{bmatrix}1&0&0&1&0&0\\ 0&-1&0&0&1&0\\ 0&0&1&0&0&1\end{bmatrix},\quad\beta_{1}^{T}=[0,7.5,0,0,-7.5,0]$
The modified ReLU function is defined as:
$\text{ReLU}(x)=\begin{cases}0&\text{if }x\leq 0\\ 1&\text{if }x>0\end{cases}$
With the input $\vec{x}_{\text{token}}^{T}=[0,6,1]$, we calculate:
$a^{1}=\text{ReLU}(\vec{x}_{\text{token}}\cdot W_{1}+\beta_{1})=[0,1,1,0,0,1]$
The vector $a^{1}$ processes the inputs for XOR and AND operations, where the middle elements act as switches to activate the respective operations. We require another layer defined by $W_{2}$ and $\beta_{2}$:
$W_{2}=\begin{bmatrix}1&0&0&0\\ 1&1&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&1&1\\ 0&0&0&1\end{bmatrix},\quad\beta_{2}^{T}=[-1.5,-1.5,-1.5,-1.5]$
$a^{2}=\text{ReLU}(a^{1}\cdot W_{2}+\beta_{2})=[0,1,0,0]^{T}$
The outputs of $a^{2}$ direct the inputs to the XOR and AND gates, assuming that $XOR$, ANDAND, and $OR$ are existing neural networks. The weights $W_{1},\beta_{1},W_{2},\beta_{2}$ used in the above example should work with all the other inputs, solving this time-dependent logical operator function.
$y=\text{OR}\Big{(}\text{XOR}(a_{2}\cdot\begin{bmatrix}1&0\\ 0&1\\ 0&0\\ 0&0\end{bmatrix}),\text{AND}(a_{2}\cdot\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&1\end{bmatrix})\Big{)}=1$
In contrast, the summation strategy adds time and concept embeddings before inputting them into the model, resulting in inputs like $\vec{x}_{\text{sum}}=[x_{1}+t_{1},x_{2}+t_{2}]$. This blending can collapse different input combinations into the same value. For example, the group $x_{1}=0$, $x_{2}=1$, $t_{1}=0$, $t_{2}=6$ would result in the same value as another group $x_{1}=0$, $x_{2}=0$, $t_{1}=0$, $t_{2}=7$, even though these inputs should produce different outputs. Once time and concept embeddings are summed, it becomes impossible to recover the original values, making it difficult for neural networks to learn time-sensitive logic. This limitation arises from the rigid functional form imposed by the summation strategy, which constrains the model’s flexibility in representing and reasoning over time intervals While such exact collisions are rare in high-dimensional space, this example highlights a structural limitation of the summation approach that can impair learning in temporally complex tasks.
We further support this argument with a simulation study, presented in Supplementary Section 9.1, which demonstrates that time tokens more effectively preserve time-sensitive logic across varying event intervals compared to the summation strategy.
3.6 Downstream Prediction Tasks
We implemented the cohorts listed in Table 2 to support evaluations of synthetic data generation and feature representation. However, two cohorts—Discharge home death and T2DM HF—were excluded from the synthetic data evaluation because they require information from the death and measurement domains, which are not included in the current version of our synthetic dataset.
Cohort
Description
Size
Median Age
Female (%)
Outcome (%)
HF readmission 30-day all-cause readmission in heart failure (HF) patients. 138,337 71 49.80 25.7
Discharge home death Mortality within 1 year since discharge to home. 227,256 50 66.89 4.41
T2DM HF Lifetime heart failure (HF) prediction since the initial diagnosis of type 2 diabetes mellitus (T2DM). 130,169 61 50.72 9.96
Hospitalization 2-year risk of hospitalization starting from the 3rd year since the initial entry into the EHR system. 722,295 45 61.99 5.91
AFib ischemic stroke 1-year risk of ischemic stroke within one year of initial diagnosis of Atrial fibrillation (AFib). 48,314 71 48.36 3.50
CAD CABG Patients with an initial diagnosis of Coronary Artery Disease (CAD) received Coronary Artery Bypass Grafting (CABG) surgery within a year, excluding prior stent grafts. 78,990 69 45.6 4.27
Table 2: Definitions and cohort characteristics of prediction tasks
3.7 Evaluation Setup
We evaluated CEHR-XGPT for three capabilities: feature representation, zero-shot prediction, and synthetic data generation. To benchmark our model against existing approaches, we compare it with MOTOR, a SOTA EHR foundation model known for its strong performance in representation learning tasks [8]. However, since MOTOR does not support zero-shot prediction or synthetic data generation, our comparison is limited to representation performance. Furthermore, we introduce an ablation baseline—GPT—which shares the same architecture as CEHR-XGPT but excludes the proposed TTE and TD learning objectives. This allows us to assess the contribution of our novel timeline design.
3.7.1 Feature Representation
We evaluated the feature representation capabilities of the models using a linear probing strategy, as proposed by Steinberg et al. and Wornow et al. [8, 5]. This approach involved extracting patient-level feature vectors from the final layer of the transformer decoder, which were then used to train a logistic regression classifier. These experiments were conducted on the disease prediction tasks described in Table 2. For comparison, we included GPT and MOTOR, both pre-trained on the Columbia dataset. The same linear probing methodology was applied to these models. Additionally, we included a logistic regression baseline trained on bag-of-words (BOW) features, as described in Section 4.3.
Beyond linear probing, we also fine-tuned CEHR-XGPT and GPT for each prediction task. The training data was split into training and validation subsets using a 90:10 ratio, with the validation set used for early stopping to prevent overfitting. We performed hyperparameter tuning using 10% of the training and validation data, running 10 trials with hyperparameters sampled uniformly: learning rates ranged from $1\times 10^{-5}$ to $5\times 10^{-5}$, and weight decay values ranged from 0.01 to 0.05. The optimal hyperparameters were selected based on the lowest validation loss. The models were then retrained on the full training set, using early stopping with a patience of 1 to monitor validation loss. MOTOR was not fine-tuned, as Steinberg et al. [8] reported that its fine-tuned and linearly probed versions exhibit comparable performance. Final performance was evaluated on the held-out test set, and we report both AUROC and AUPRC metrics. The 95% confidence intervals were computed using bootstrapping.
3.7.2 Zero-Shot Prediction
Our evaluation followed the zero-shot setup proposed by ETHOS [12], in which the patient sequence up to the prediction time was provided to CEHR-XGPT to generate 50 simulated patient trajectories. Generation was terminated when either of the following conditions was met: (1) the generated tokens exceeded the defined prediction window, or (2) the outcome event(s) occurred within the prediction time. If the [END] token was encountered within the prediction window—causing premature termination—we treated the simulation as censored and discarded it, replacing it with a new one. The probability of the predicted outcome was then estimated as the proportion of simulations in which the event occurred. Using these estimated probabilities and the corresponding ground-truth labels, we computed AUROC and AUPRC scores. The standard deviation of these estimates was assessed via bootstrapping. An illustrative example of this zero-shot prediction setup is provided in Supplementary Section 9.4. Due to the computational cost associated with the generative nature of zero-shot inference, we randomly sampled 5,000 test cases from each prediction task.
3.7.3 Synthetic Data Generation
We extended the synthetic data generation procedure developed by Pang et al. [35], addressing one of the primary challenges identified in their work: optimizing the sampling strategy. To this end, we adopted a mixture of experts approach, consisting of the following steps:
  1. 1.
    Data Generation: Multiple synthetic datasets were generated using different combinations of sample sizes, model checkpoints, generation temperatures, top_p, top_k, and repetition_penalty values.
  2. 2.
    Data Pooling and Conversion: All synthetic patient sequences were pooled and converted into a unified OMOP instance.
The resulting synthetic OMOP dataset contained approximately 3.7 million patient records, and details of this synthetic dataset can be found in Supplementary Section 9.5. Based on this dataset, we constructed four prediction tasks using methodologies described in [11, 35, 39]. Cohort definitions are summarized in Table 2.
We evaluated the utility of the synthetic data under two real-world simulation scenarios:
  1. 1.
    Model Development: Training models solely on synthetic data.
  2. 2.
    Training Set Augmentation: Augmenting real training data with synthetic data. In both cases, performance was measured on a real-world test set.
Feature extraction followed a bag-of-words (BOW) approach, counting concept frequency within defined observation windows. Logistic regression models were trained using Scikit-Learn’s default settings. For real data baselines, we used the original train/test splits.
In both cases, performance was measured on a real-world test set. Feature extraction followed a bag-of-words (BOW) approach, counting concept frequency within defined observation windows. Logistic regression models were trained using Scikit-Learn’s default settings. For real data baselines, we used the original train/test splits. The standard deviation was estimated through bootstrapping.
To assess the fidelity of the synthetic dataset in capturing complex longitudinal patterns, we replicated the treatment pathway analyses for Hypertension (HTN), Diabetes, and Depression originally proposed by Hripcsak et al. [40]. These analyses impose strict inclusion criteria, such as requiring a three-year follow-up period for HTN patients with continuous exposure to antihypertensive medications across nine consecutive 120-day intervals. As a result, the final cohorts are typically small and highly selective, making them a sensitive test of temporal coherence. The full study design is illustrated in Figure 3.
Figure 3: Treatment pathway study cohort criteria: Patients were required to have at least 365 days of observation prior to their first drug exposure for the disease of interest. Following this initial exposure, patients had to maintain continuous treatment, defined as having at least one drug exposure every 120 days over a 3-year period.
Finally, we evaluated the privacy risks associated with synthetic data sharing using the framework introduced by Yan et al. [41]. This framework includes four types of attacks: Membership Inference [42], Attribute Inference [23], Meaningful Identity Disclosure [43], and Nearest Neighbor Adversary Attacks (NNAA) [44]. We adopted the evaluation metrics and decision thresholds directly from Yan et al., and provide additional implementation details in Supplementary Section 9.5.4.
3.7.4 Generalization on EHRShot Benchmark
To evaluate the generalizability of CEHR-XGPT, we utilized the ehrshot benchmarking dataset, which includes 6,739 comprehensive de-identified EHR records from Stanford Medicine [9]. This dataset is structured into 15 manually created tasks—both binary and multi-class classification tasks—across three categories: Operational Outcomes (e.g., 30-day readmission), Assignment of New Diagnoses (e.g., 1-year risk of pancreatic cancer), and Anticipating Lab Test Results. Given that CEHR-XGPT was specifically pre-trained on visits, conditions, procedures, and medications, we excluded the Anticipating Lab Test Results category from our analysis.
We opted for the OMOP version of ehrshot for our experiments as both CEHR-XGPT and GPT were trained using the CUIMC OMOP, where medical events are coded using the standard OMOP vocabularies. For each task, we extracted all relevant events prior to the prediction time from the ehrshot OMOP and transformed them into the CEHR-XGPT patient representation. We evaluated both linear probing and full fine-tuning on the downstream tasks, similar to the procedure described in Section 4.1. In both settings, we extracted the vector corresponding to the final token of the patient sequence from the CEHR-XGPT output and used it to train a linear classifier. The key distinction between the two approaches is that in linear probing, the pre-trained model weights were frozen, whereas in full fine-tuning, all weights were updated during training. For linear probing, we used the original tokenizer vocabulary and excluded any events not present in it. For fine-tuning, we expanded the tokenizer vocabulary to include over 5000 standard OMOP concept IDs specific to the Stanford EHR data, which were absent in the Columbia EHR. We conducted a hyperparameter search over 10 trials, sampling learning rates uniformly between $1\times 10^{-5}$ and $1\times 10^{-4}$. Sample packing was used during training, with a maximum of 16,384 tokens allowed per batch.
For model comparisons, we included all models listed on the ehrshot leaderboard. Furthermore, we considered the best-performing model from a recent study by the authors of Context Clues, which evaluated several causal language model architectures (GPT2, Llama, Mamba, etc.) with various context windows on Stanford EHR data [5]. Their findings, reported on the ehrshot, served as an additional baseline in our experiments.
4 Experiments and Results
Our model, CEHR-XGPT, demonstrates versatility across three key capabilities: feature representation, zero-shot prediction, and synthetic data generation. For comparison, we include MOTOR as a SOTA EHR foundation model focused on representation learning. Additionally, we evaluate a variant of CEHR-XGPT, referred to as GPT, which excludes TD and TTE learning objectives.
GPT is assessed under the same experimental conditions as CEHR-XGPT for the three aforementioned capabilities, except the large-scale synthetic data generation experiment, which was excluded due to the computational cost of generating 3.7 million synthetic patients. Instead, we evaluate its long-term generation ability through a targeted analysis of the treatment pathway using a smaller synthetic cohort of 100,000 patients.
Finally, we assess the generalizability of CEHR-XGPT using the ehrshot benchmarking dataset [9], which provides a diverse range of clinical prediction tasks.
4.1 Feature Representation
To evaluate CEHR-XGPT’s ability to generate meaningful patient representations, we assessed its performance under both linear probing and fine-tuning settings, and compared it with MOTOR and GPT. The Receiver Operating Characteristic (ROC) curves are shown in Figure 4, while the Precision-Recall (PR) curves are provided in Supplementary Figure 6. We reported ROC-AUC and PR-AUC across all tasks in Table 3.
Under the linear probing setting, MOTOR achieved the highest AUROC and AUPRC across all tasks, outperforming both CEHR-XGPT-L and GPT-L. This may reflect differences in pretraining strategies, as MOTOR is explicitly optimized for representation learning (e.g., time-to-event prediction), whereas CEHR-XGPT is trained using next-token prediction.
With full model fine-tuning, both CEHR-XGPT-T and GPT-T showed substantial improvements across all tasks compared to their linear probing counterparts. Both models also outperformed MOTOR on several tasks, including HF Readmission, Hospitalization, AFib Ischemic Stroke, and CAD CABG. The largest gains were observed on the AFib Ischemic Stroke task, where fine-tuning improved AUROC by 7% and AUPRC by 45%. While MOTOR demonstrated minimal performance gains from fine-tuning in prior work [8], CEHR-XGPT showed competitive results, particularly when fine-tuned.
Comparing GPT and CEHR-XGPT, we observed that the incorporation of time token–specific learning objectives had minimal impact on supervised prediction tasks (Table 3). However, as described in the next sections, these objectives showed its clear benefits for zero-shot generalization and synthetic data generation.
Figure 4: Receiver Operating Characteristic (ROC) curves comparing the performance across six clinical prediction tasks. Models are distinguished by unique color and line style combinations. Performance metrics show Area Under the Curve (AUC) values with 95% confidence intervals calculated using bootstrap sampling. Models are ranked in descending order of AUC performance within each subplot, with annotations positioned in the bottom right corner. The shared legend displays all evaluated models with their corresponding visual representations. Abbreviations: CEHR-XGPT-L (CEHR-XGPT with linear probing), CEHR-XGPT-T (CEHR-XGPT with fine-tuning), GPT-L (The CEHR-XGPT original variant with linear probing), GPT-T (The CEHR-XGPT original variant with fine-tuning), LR Baseline (Logistic Regression).
Cohort Metric
LR
Baseline
CEHR
GPT-L
GPT-L
CEHR
GPT-T
GPT-T
MOTOR
HF readmission
AUROC
AUPRC
64.1 $\pm$ 0.3%
35.1 $\pm$ 0.2%
68.0 $\pm$ 0.3%
37.4 $\pm$ 0.5%
67.9 $\pm$ 0.3%
37.6 $\pm$ 0.5%
70.1 $\pm$ 0.3%
40.5 $\pm$ 0.5%
69.5 $\pm$ 0.3%
39.8 $\pm$ 0.5%
69.2 $\pm$ 0.3%
39.2 $\pm$ 0.5%
T2DM HF
AUROC
AUPRC
77.9 $\pm$ 0.7%
28.3 $\pm$ 1.3%
83.1 $\pm$ 0.3%
38.8 $\pm$ 0.8%
83.1 $\pm$ 0.3%
38.3 $\pm$ 0.8%
83.5 $\pm$ 0.3%
39.2 $\pm$ 0.8%
83.3 $\pm$ 0.3%
38.2 $\pm$ 0.8%
85.0 $\pm$ 0.3%
40.2 $\pm$ 0.8%
Hospitalization
AUROC
AUPRC
76.2 $\pm$ 0.2%
20.3 $\pm$ 0.3%
82.5 $\pm$ 0.2%
31.7 $\pm$ 0.4%
82.2 $\pm$ 0.2%
30.1 $\pm$ 0.4%
85.1 $\pm$ 0.2%
39.2 $\pm$ 0.4%
85.2 $\pm$ 0.2%
38.9 $\pm$ 0.5%
85.0 $\pm$ 0.2%
38.4 $\pm$ 0.5%
Discharge mortality
AUROC
AUPRC
93.7 $\pm$ 0.3%
51.0 $\pm$ 1.7%
94.6 $\pm$ 0.1%
53.6 $\pm$ 0.8%
94.6 $\pm$ 0.1%
53.4 $\pm$ 0.8%
95.0 $\pm$ 0.1%
56.6 $\pm$ 0.8%
94.9 $\pm$ 0.1%
55.5 $\pm$ 0.8%
96.6 $\pm$ 0.1%
70.7 $\pm$ 0.7%
AFib ischemic stroke
AUROC
AUPRC
70.5 $\pm$ 1.1%
9.0 $\pm$ 0.7%
74.1 $\pm$ 0.9%
11.1 $\pm$ 0.9%
73.3 $\pm$ 0.9%
10.6 $\pm$ 0.9%
78.3 $\pm$ 0.8%
14.9 $\pm$ 1.3%
77.5 $\pm$ 0.9%
13.7 $\pm$ 1.1%
76.3 $\pm$ 0.9%
17.3 $\pm$ 1.4%
CAD CABG
AUROC
AUPRC
84.8 $\pm$ 0.7%
33.3 $\pm$ 1.6%
91.0 $\pm$ 0.4%
53.7 $\pm$ 1.4%
90.6 $\pm$ 0.5%
52.5 $\pm$ 1.4%
91.7 $\pm$ 0.4%
56.9 $\pm$ 1.4%
91.3 $\pm$ 0.4%
57.1 $\pm$ 1.4%
90.8 $\pm$ 0.5%
54.6 $\pm$ 1.3%
Table 3: Model performance on each disease prediction task. Abbreviations: CEHR-XGPT-L (CEHR-XGPT with linear probing), CEHR-XGPT-T (CEHR-XGPT with fine-tuning), GPT-L (The CEHR-XGPT original variant with linear probing), GPT-T (The CEHR-XGPT original variant with fine-tuning), LR Baseline (Logistic Regression).
4.2 Zero-Shot Prediction
We evaluated the zero-shot prediction capability of CEHR-XGPT using three cohorts described in Table 2, representing short-term, mid-term, and long-term risk prediction tasks: HF Readmission (30-day risk), CAD CABG (1-year risk), and T2DM HF (lifetime risk).
As shown in Table 4, CEHR-XGPT achieved strong zero-shot performance across all three tasks. On the short-term task (HF Readmission), it performed comparably to GPT, while demonstrating slightly improved performance on the mid-term and long-term tasks (CAD CABG and T2DM HF, respectively). These results highlight the benefits of CEHR-XGPT’s temporal modeling capabilities enabled by time tokens and its autoregressive design, which are especially important in longer-horizon prediction. Importantly, the zero-shot performance of CEHR-XGPT (without task-specific fine-tuning) not only rivals the linear probing performance of pretrained models, but also outperforms count-based baselines on both HF Readmission and CAD CABG, and matches the best baseline on T2DM HF. This underscores CEHR-XGPT’s ability to generalize to diverse clinical prediction tasks.
Cohort
CEHR-XGPT
GPT
HF readmission
(30 day risk)
R = 66.8 $\pm$ 0.8%
P = 35.1 $\pm$ 1.4%
R = 66.8 $\pm$ 0.9%
P = 35.8 $\pm$ 1.4%
T2DM HF
(lifetime risk)
R = 78.9 $\pm$ 0.9%
P = 27.3 $\pm$ 1.6%
R = 78.7 $\pm$ 0.9%
P = 26.9 $\pm$ 1.5%
CAD CABG
(1-year risk)
R = 90.9 $\pm$ 1.1%
P = 55.6 $\pm$ 3.3%
R = 90.0 $\pm$ 1.2%
P = 54.9 $\pm$ 3.4%
Table 4: Zero-shot prediction performance on the short-term prediction task HF Readmission, mid-term prediction task CAD CABG, long-term prediction task T2DM HF. Abbreviations: R (AUROC), P (AUPRC)
4.3 Synthetic Data Generation
Table 5 presents model performance results using synthetic data. For two tasks—Hospitalization and AFib Ischemic Stroke—models trained entirely on synthetic data performed comparably to those trained on real data when evaluated on the real test set. In contrast, performance for HF Readmission and CAD CABG was slightly lower with synthetic-only training. However, augmenting the real training set with synthetic data led to modest yet consistent improvements in test performance across all four tasks, relative to models trained on real data alone, suggesting that CEHR-XGPT–generated data can serve as a useful supplement when real data is limited.
Cohort
M
Real
Syn
Syn + Real
HF
Readmit
R
P
64.1 $\pm$ 0.3%
35.1 $\pm$ 0.2%
[…]
32.0 $\pm$ 0.4%
64.5 $\pm$ 0.3%
34.5 $\pm$ 0.5%
Hospital-
ization
R
P
76.2 $\pm$ 0.2%
20.3 $\pm$ 0.3%
76.2 $\pm$ 0.2%
22.3 $\pm$ 0.4%
77.7 $\pm$ 0.2%
23.8 $\pm$ 0.4 %
AFib IS
R
P
70.5 $\pm$ 1.1%
9.0 $\pm$ 0.7%
70.7 $\pm$ 1.1%
9.7 $\pm$ 0.9%
71.2 $\pm$ 1.1%
10.4 $\pm$ 0.9%
CAD
CABG
R
P
84.8 $\pm$ 0.7%
33.3 $\pm$ 1.6%
82.9 $\pm$ 0.7%
26.5 $\pm$ 1.5%
85.0 $\pm$ 0.7%
32.6 $\pm$ 1.6%
Table 5: Logistic regression model performance (AUROC and AUPRC) across different datasets using synthetic datasets generated from CEHR-XGPT. The columns compare the model performance when 1) trained on synthetic data alone and tested on the real test set, and 2) trained on synthetic + real data, tested on the real test set. Abbreviations: HF Readmit(HF Readmission), AFib IS (AFib Ischemic Stroke), M(Metric), R (AUROC), P (AUPRC)
In treatment pathway analyses, the synthetic dataset generated by CEHR-XGPT closely matched real-world prevalence for HTN and Depression, though it underestimated Diabetes (Supplementary Table 10): HTN 0.41% vs. 0.45%, Diabetes 0.11% vs. 0.18%, and Depression 0.18% vs. 0.14%. In contrast, the GPT dataset underestimated HTN (0.26%) and Diabetes (0.10%) while only aligning on Depression (0.14%). These findings underscore the value of timeline-aware modeling, with CEHR-XGPT more faithfully preserving long-term patient trajectories than GPT.
Privacy evaluation results are shown in Supplementary Table 11. Across all four attack types, CEHR-XGPT’s synthetic dataset achieved risk scores below 0.333, a threshold recommended by Yan et al. to indicate minimal privacy risk. These results demonstrate that CEHR-XGPT can generate synthetic EHR data with low measurable privacy risk while preserving clinical utility for downstream modeling.
4.4 Generalization on EHR-SHOT Benchmark
We demonstrated generalizability of CEHR-XGPT by evaluating it on the external ehrshot benchmarking dataset, which enables comparison with several competitive models available on the ehrshot Leaderboard, including MOTOR, Context, CLMBR, GBM, LR, RF, and GPT (Table 6 and Supplementary table 12). Overall, CEHR-XGPT outperformed GPT in both fine-tuned and linear-probing settings. For Patient Outcomes, fine-tuning CEHR-XGPT improved performance, with average gains of 1.2% in AUROC and 5.0% in AUPRC compared to linear probing. In contrast, on New Diagnosis tasks, fine-tuning CEHR-XGPT yielded a higher AUROC with an average gain of 1.9% but slightly reduced AUPRC by 0.6%. Interestingly, GPT followed a different pattern: for Patient Outcomes, fine-tuning and linear probing performed similarly, with fine-tuning achieving a modest average improvement of 0.9% in AUPRC. However, for New Diagnosis, fine-tuning degraded performance relative to linear probing, particularly for two low-prevalence cohorts, Celiac and Pancreatic Cancer. A similar trend was observed for CEHR-XGPT on Pancreatic Cancer, where linear probing outperformed fine-tuning. These results suggest that under extreme class imbalance, linear probing may offer a more robust approach for predicting rare events than fine-tuning.
More importantly, both fine-tuned and linear-probing versions of CEHR-XGPT demonstrated strong performance across these tasks, achieving the best results on New Diagnosis prediction compared to all other models. Notably, CEHR-XGPT outperformed models such as CLMBR, despite the latter being trained on data drawn from the same distribution (Stanford EHR data) as the ehrshot dataset. This suggests that simply extending the context window or adopting alternative architectures does not necessarily yield better performance on these tasks, and that CEHR-XGPT’s novel patient representation design and time-specific learning objectives may be key contributors to its improved performance.
Task Prev Metric MOTOR Context Best Baseline CEHR GPT-T CEHR GPT-L GPT-T GPT-L
Patient
Outcome
R
P
83.6%
49.0%
83.3%
N/A
82.4%
43.7%
83.3%
46.7%
82.1%
41.7%
82.6%
44.0%
81.7%
40.8%
ICU
Admission
4.5%
R
P
84.8%
32.4%
[…]
Long LOS
25.2%
R
P
81.4%
57.6%
[…]
30-day
Readmission
13.0%
R
P
81.0%
41.2%
[…]
New
Diagnosis
R
P
75.6%
19.3%
75.6%
N/A
74.9%
21.2%
75.9%
20.8%
74.0%
21.4%
72.0%
20.4%
72.8%
21.5%
Acute MI
6.8%
R
P
74.7%
18.4%
[…]
Celiac
1.3%
R
P
75.8%
23.1%
[…]
Pancreatic
Cancer
3.8%
R
P
88.5%
47.2%
[…]
Hypertension
13.7%
R
P
71.8%
25.8%
[…]
Lupus
2.2%
R
P
79.3%
17.9%
[…]
Hyperlipidemia
12.7%
R
P
72.0%
25.8%
[…]
Table 6: Performance metrics of selected models for patient outcome and new diagnosis predictions using the ehrshot dataset. This table summarizes both individual and average model performance across tasks. The Best Baseline column shows the best metric among CLMBR, GBM, LR, and RF. For MOTOR, only aggregated performance data is available since specific task details were not published on the ehrshot leaderboard. For entries marked as ’Context,’ only the highest AUROC is reported due to the absence of AUPRC metrics in the original publication. Abbreviations: CEHR-XGPT-L (CEHR-XGPT with linear probing), CEHR-XGPT-T (CEHR-XGPT with fine-tuning), GPT-L (The CEHR-XGPT original variant with linear probing), GPT-T (The CEHR-XGPT original variant with fine-tuning). Prev (Prevalence), R (AUROC), P (AUPRC)
5 Discussion
CEHR-XGPT is, to our knowledge, the first EHR foundation model comprehensively evaluated across three core capabilities: feature representation, zero-shot prediction, and synthetic data generation. While prior models typically specialize in one area, CEHR-XGPT demonstrates broad versatility, enabling a wide range of clinical tasks without requiring task-specific retraining. This flexibility is particularly valuable in real-world settings with limited labeled data, evolving populations, or rapidly changing outcome definitions.
A key enabler of this versatility is CEHR-XGPT’s use of time tokens, along with Time Decomposition (TD) and Time-to-Event (TTE) objectives. These mechanisms help encode temporal structure directly into patient sequences and appear to enhance generalizability—particularly in zero-shot and transfer settings. For example, in the external ehrshot benchmark, CEHR-XGPT achieved strong performance through vocabulary expansion and full fine-tuning. It matched top-reported models on Patient Outcome tasks and outperformed all previously reported models on New Diagnosis tasks. While MOTOR, CLMBR, Llama, and Mamba were evaluated using linear probing in prior studies, their performance could improve with fine-tuning.
In contrast to MOTOR, which consistently outperformed other models in linear probing, CEHR-XGPT’s representation performance was more modest. This may reflect differences in pretraining: MOTOR was trained to forecast thousands of future events across time intervals, promoting long-range structure learning. CEHR-XGPT, by comparison, was trained with next-token prediction and auxiliary temporal objectives—favoring sequential modeling over embedding quality. These results suggest that the alignment between pretraining and downstream tasks plays a critical role in determining the effectiveness of linear probing. Another factor is feature coverage: CEHR-XGPT currently excludes measurements and observations, domains used by MOTOR. Expanding CEHR-XGPT’s input domains may help close this gap.
Interestingly, GPT (CEHR-XGPT without time-specific learning objectives) demonstrated similar performance to CEHR-XGPT on in-distribution representation tasks, but underperformed in zero-shot prediction, treatment pathway analysis using synthetic data, and external validation. This suggests that time tokens act as a form of structural regularization—limiting overfitting to the training distribution while improving temporal generalization. Analogous to regularization in traditional models, the TD and TTE objectives guide the model to learn more transferable temporal representations, which is critical for generalization across health systems.
It is also worth emphasizing that feature representation and zero-shot prediction serve different clinical purposes. Representation learning supports embedding-based tasks like clustering or risk stratification, while zero-shot prediction enables outcome forecasting without retraining. In settings where labeled data or retraining resources are scarce—such as low-resource hospitals, emerging diseases, or real-time triage—zero-shot capabilities may offer faster and more scalable deployment.
In the context of synthetic data generation, most prior frameworks—such as Theodorou et al. [6]—adopt a conditional generation strategy to synthesize patients based on predefined conditions like T2DM. However, this approach relies on condition-specific labels assigned in advance and does not generalize well to novel or unseen conditions. In contrast, our method models the full sequence distribution of the Columbia patient population using a specialized representation that preserves complete patient timelines. A key distinction of our approach is its ability to convert generated sequences back into a common data model, enabling researchers to identify patients with any condition post hoc using standard tools such as pandas or SQL.
Taken together, these results position CEHR-XGPT as a flexible and general-purpose foundation model for EHR data. Its support for zero-shot prediction, synthetic data generation, and representation learning makes it well suited for diverse clinical workflows, including cohort discovery, surveillance, and rapid model prototyping. Future work may extend CEHR-XGPT by incorporating additional domains (e.g., measurements, labs, observations), improving computational efficiency (e.g., via adapters or distillation), and enhancing clinical interpretability through human-in-the-loop evaluation.
6 Data availability
Due to institutional policies regarding the sharing of patient-level EHR data, we are unable to share the dataset used for training. However, all code, along with instructions to replicate our work, is accessible on GitHub at https://github.com/knatarajan-lab/cehrgpt. This will enable institutions with an OMOP database to train models and generate synthetic data within their internal environment.
7 Acknowledgement
This work was supported by the National Institutes of Health (5U2COD023196, 3OT2OD026556, 5U01TR002062, 5R35GM147004). The funding sources supported individuals in curating and analyzing electronic health record data, which spurred the idea of this work.
8 Author contributions
CP conceived the study. CP led the conceptualization and methodology design with contributions from XJ, CP. CP and XJ conducted the data curation and cohort design. CP, XJ carried out the formal analysis, including various evaluation metrics. CP led the project administration. CP, JP, and XJ authored the original draft, which was reviewed and edited by KN, SJ, and NE. All authors had final responsibility for the decision to submit for publication.